### v1

# TAROT: Task-Oriented Authorship Obfuscation Using Policy Optimization Methods

Abstract
Authorship obfuscation aims to disguise the identity of an author within a text by altering the writing style, vocabulary, syntax, and other linguistic features associated with the text author. This alteration needs to balance privacy and utility. While strong obfuscation techniques can effectively hide the author’s identity, they often degrade the quality and usefulness of the text for its intended purpose. Conversely, maintaining high utility tends to provide insufficient privacy, making it easier for an adversary to de-anonymize the author. Thus, achieving an optimal trade-off between these two conflicting objectives is crucial. In this paper, we propose TAROT: Task-Oriented Authorship Obfuscation Using Policy Optimization, a new unsupervised authorship obfuscation method whose goal is to optimize the privacy-utility trade-off by regenerating the entire text considering its downstream utility. Our approach leverages policy optimization as a fine-tuning paradigm over small language models in order to rewrite texts by preserving author identity and downstream task utility. We show that our approach largely reduce the accuracy of attackers while preserving utility. We make our code and models publicly available. [^1] https://github.com/hornetsecurity/tarot
1 Introduction
Privacy is a critical consideration when developing and deploying trustworthy machine learning systems. With the expansion of large language models (LLM), large datasets are used for pretraining and fine-tuning models on a wide variety of tasks. With this comes a risk of sensitive or personal information being exposed or misused. Text anonymization is a technique that can address these privacy concerns. It involves the process of removing or obfuscating the author’s personal information from textual data. This helps protect the privacy of individuals whose data is used to train models, however, it needs to be performed in a way that still allows models to learn useful patterns and insights from texts.
Figure 1: Illustration of the two versions of TAROT: We generate obfuscation candidates and optimize the best policy using Reinforcement Learning and preference optimization.
Currently, most work done on text anonymization focuses on redacting sensitive entities in a given document Lison et al. (2021). This is sufficient for texts where the only private aspects are entities, such as medical reports, court cases, or biographies. But it is inadequate for removing author’s writing style, or weak signals that can be used as hints for identification, which is for example the case for blog articles or emails. Redacting entities in text while keeping stylometric features linked to a specific individual would eventually result in a leak of information. Indeed, the writing style is a strong indicator of a person’s identity Mosteller and Wallace (1963). Prior work done on authorship attribution highlights the large quantity of information that can be extracted from seemingly anonymized texts and the ease of identification of authors, especially for long documents Fabien et al. (2020).
To solve this issue, authorship obfuscation (AO) aims to hide the author’s identity by replacing some part of the text associated with authorship indicators. Modifying the original text can impact its usability for specific tasks (i.e. utility), and therefore badly affects the downstream performances and text comprehension of machine learning models. The enforcement of privacy creates a trade-off between privacy and utility, where keeping the original text preserves the unchanged utility of the text, while not defending against attribution attacks. On the other side, obfuscating the entire text guarantees privacy, but leads to unusable text in practice.
Previous approaches design their obfuscation by maximizing the preserved text content. They limit the modifications to small and targeted edits in order to preserve text meaning and keep textual content as close as possible to the original. As a result, those approaches often lead to insufficient modification in the text, especially against realistic attack scenarios Zhai et al. (2022).
To address these limitations, we reframe the AO problem into an adversarial problem between two adversaries (e.g. machine learning models): one attacker model whose goal is to reveal the identity of a given author from written texts, and one utility model that aims to perform a given task using authors data. The goal is to provide a modified version of the original text such that the utility model can accurately perform its task while preventing the attacker from identifying the author, making the obfuscation task-oriented. The notion of task-oriented obfuscation/anonymization also takes its origin in the law. As stated by GDPR European Parliament and Council of the European Union (2016), the collection and process of personal information (including written texts) have to be specified for a given usage.
In order to learn this privacy-utility trade-off, we use the combination of supervised fine-tuning (SFT) and policy optimization (PO) to guide a generative model into generating privacy- and utility-preserving outputs. Our model learns to rewrite the text while removing potential authorship signals, and preserving the text utility for a downstream task. This rewriting goal is further validated by the conclusion of Weitzenboeck et al. (2022) which showed how difficult it is to comply with GDPR requirements concerning text anonymization without changing the entire text.
We fine-tune a text simplification model for AO using a customized reward model. We design an unsupervised reward model for PO using 2 pretrained sentence embedding models. The utility reward, penalizes the fact that the General Text Embeddings Li et al. (2023) of the anonymized sentence is too far removed from that of the original sentence. The author rewards does the opposite on the embedding built by the Universal Authorship Representation model from Rivera-Soto et al. (2021). Our final models are trained in an open-world setting where the number of authors is not defined, same goes for the end utility for our model to work on a multi-task setting. We also provide experimentation on three different datasets, movie reviews, blog articles and scholar documents. We show that TAROT can be used on multiple datasets targeting different tasks while protecting authorship.
In summary, we list the main contributions as follows:
  • We design a new framework for task-oriented AO by leveraging PO algorithms to maximize the end usage of data. The objective is to help reduce the traditional constraints associated with utility preservation in the literature (strict content preservation and semantic quality) by looking for a downstream classification task to achieve with the anonymized data.
  • Starting from this framework, we propose TAROT, a task-oriented generation model aiming to obfuscate text without any prior knowledge of the author (making it unsupervised, and usable on any dataset, even if the authors are not clearly indicated) while maximizing the utility for a variety of tasks. We release two versions of TAROT from two different fine-tuning PO algorithms: TAROT-PPO and TAROT-DPO.
  • We further evaluate TAROT on three datasets associated with different classification tasks, using different authorship attackers and downstream usage scenarios.
2 Related Work
Authorship Obfuscation
Obfuscation techniques can be regrouped into two categories, depending on their implementation. Generic methods, on one hand, are methods that were not explicitly designed for AO, but show interesting performance. These methods include machine translation Altakrori et al. (2022); Keswani et al. (2016), paraphrasing Krishna et al. (2023), or synonym replacements Potthast et al. (2016).
More recently, advanced techniques were built explicitly for AO, often relying on a trained attacker performing authorship attribution attacks on the obfuscated text. Then, they perform accurate adversarial text edits from the attacker knowledge on authors in order to obtain a privatized output. Mutant-X Mahmood et al. (2019), is a genetic algorithm that utilizes GloVE Pennington et al. (2014) word embeddings selected from an SVM or Random Forest attacker to replace words in a document with similar ones.
Jamdec Fisher et al. (2024) is an unsupervised approach for obfuscating the writing style of text while preserving semantics. It uses embedding-based and likelihood-based methods, rather than attacker-based methods, to extract keywords, then generates multiple text variations using Constrained Diverse Beam Search on GPT2-XL (1.61B parameters). Finally, the candidates are filtered using Natural Language Inference (NLI) and Corpus of Linguistic Acceptability (CoLA) metrics to ensure coherence, content preservation, and grammatical correctness.
Recently, ALISON Xing et al. (2024) employs a lightweight multilayer perceptron classifier using part-of-speech (POS) sequences to guide obfuscation, and leverages a BERT pre-trained language model to generate replacement sequences. By ranking and replacing important POS n-grams, ALISON obfuscates text uniformly, reducing classifier confidence.
Reinforcement Learning.
In NLP, reinforcement learning (RL) is often used to capture small signals over word or sentence embedding. For example, Mosallanezhad et al. (2019) proposes a text representation anonymization approach that employs deep reinforcement learning to detect and modify text embeddings to maintain a good privacy-utility trade-off.
With the expansion of Reinforcement Learning from Human Feedback (RLHF) as a LLM fine-tuning paradigm, RL techniques have been leveraged to improve language models with scalar metrics by optimizing rewards from (human) feedbacks. It has emerged as a prominent tool for tackling undesirable behaviors such as toxicity, social biases, and offensive language Ouyang et al. (2022). This is accomplished by implementing PO algorithms to optimize a language model (LM) by associating a reward with each generation, derived from a trained reward model.
Very recently, Liu et al. (2024) introduced an authorship style transfer method using PO. They optimize style transfer generation using style similarity reward models. Authorship style transfer is similar to AO in the way those task’s goal is to change within a text the author writing style. However, style transfer assumes a distinct target style to achieve, whereas AO assumes a lack of distinct style. Fisher et al. (2024) also showed the ineffectiveness of style transfer for AO. To the best of our knowledge, our work is the first one applying PO algorithms on AO.
3 Methodology
3.1 Problem Formulation
Let $x_{\text{ori}}$ represent the original document authored by a specific author $a\in\mathcal{A}$. $\mathcal{A}$ denoting a predetermined set of authors. The objective of authorship obfuscation is to generate a new document, denoted as $x_{\text{obf}}$, which cannot be attributed to the original author $a$. To assess the effectiveness of obfuscation, we employ a classification model, denoted as $f_{priv}(\cdot)$ (i.e. an authorship attribution model), which has been trained to distinguish documents based on their respective authors within $\mathcal{A}$. The goal of authorship obfuscation is to design an obfuscation method $O(\cdot)$, such that $f_{priv}(O(x_{\text{ori}}))\neq f_{priv}(x_{\text{ori}}).$
In addition, a successful obfuscation algorithm would not only trick an attacker into predicting the wrong author, but also preserve the document utility for downstream usage. In this paper, instead of mainly measuring this utility change though various semantic or content preservation metrics (i.e. METEOR score, BERT score, …) we highlight the selection of a prior task $\mathcal{T}$ in order to evaluate obfuscation with respect to $\mathcal{T}$. We denote as $f_{\mathcal{T}}(\cdot)$ the classification model used for a utility task. An ideal $O(\cdot)$ would preserve the original label $f_{\mathcal{T}}(O(x_{\text{ori}}))=f_{\mathcal{T}}(x_{\text{ori}})$.
Note that $\mathcal{T}$ is likely not known when we train the obfuscation model, underscoring the necessity for a versatile obfuscation strategy. This task-agnostic approach prevents the obfuscation model from learning to transform the text specifically to fit the label of $\mathcal{T}$, which would compromise its generality across different tasks.
3.2 Framework Overview
Our task-oriented framework can be decomposed in two steps. First, we initialize our generation model from a SFT baseline, this will first guide our LM to generate modified versions of the input text instead of proceeding text copy. Second, we apply a PO algorithm to fine-tune our SFT model. We experiment with two different PO algorithms, Proximal Policy Optimization Schulman et al. (2017) and Direct Preference Optimization Rafailov et al. (2023) (see Figure 1). We optimize our SFT generations using a reward model composed of both privacy and content preservation components.
3.3 Supervised fine-tuning
First, we use a fine-tuned LM to initiate our text generation task. We employ the Keep It Simple [^2] https://hf.co/philippelaban/keep_it_simple simplification model Laban et al. (2021) as an SFT baseline. This model is a fine-tuned version of GPT2-medium on the Newsela [^3] https://newsela.com/ dataset for text simplification. The utilization of a simplification model encourages a reduction in the amount of information conveyed by a sentence, thereby affording the opportunity to eliminate author-specific features. To our knowledge, this is the first time that a simplification model has been used for AO. Moreover, our framework is broadly compatible with any autoregressive LM, and can be adapted with larger architectures and other generation tasks.
Dataset Authors Texts
Avg. Texts / Author
Avg. Words / Text
Avg. Tokens / Text
Avg. Chars / Text
IMDb $10$ $01000$(±0)1000plus-or-minus01000($\pm$ 0)1000 ( ± 0 ) $364(\pm 209)$ $393(\pm 228)$ $1869(\pm 1077)$
$20$ $20000$ $1000(\pm 0)$ $345(\pm 209)$ $371(\pm 225)$ $1767(\pm 1081)$
BAC $10$ $42353$(±639)2353plus-or-minus6392353($\pm$ 639)2353 ( ± 639 ) $118(\pm 195)$ $120(\pm 236)$ $524(\pm 1027)$
$20$ $39379$ $1969(\pm 599)$ $118(\pm 175)$ $123(\pm 214)$ $529(\pm 921)$
AMT $10$ $196$ $20(\pm 2)$ $497(\pm 14)$ $592(\pm 41)$ $2956(\pm 194)$
$20$ $362$ $18(\pm 2)$ $502(\pm 102)$ $590(\pm 38)$ $2956(\pm 207)$
Table 1: Dataset statistics
3.4 Policy Optimization Algorithms
We use two different PO algorithms to optimize generations of our SFT baseline. The Proximal Policy Optimization (PPO) Schulman et al. (2017) algorithm is a policy gradient method whose goal is to optimize a policy with respect to continuous rewards. In our case, a policy is a generation strategy, i.e. a final LM. Initialized from the SFT policy, we sample completions $y$ given prompts $x$ and the reward model parametrized by $\phi$ produces a score $r_{\phi}(x,y)$ based on these completions. The reward score $r_{\phi}(x,y)$ is then combined with a Kullback–Leibler (KL) penalty to ensure the policy does not deviate too much from the SFT policy (leading to unusable generations). Specifically, the reward of the RL problem is:
$R(x,y)=r_{\phi}(x,y)-\beta\mathbb{D}_{\textrm{KL}}\bigl{[}\pi_{\theta}(y\mid x)\mid\mid\pi_{\text{SFT}}(y\mid x)\bigr{]}$
where $\beta$ is a parameter controlling the strength of the KL penalty, $\theta$ the parameters of RL policy $\pi_{\theta}$, and $r_{\phi}$ the reward model with parameters $\phi$. Then, PPO is used to maximize the following objective:
$\max_{\pi_{\theta}}\ \mathbb{E}_{x\sim\mathcal{D}_{\text{SFT}},y\sim\pi_{\theta}(y\mid x)}R(x,y)$
where $\mathcal{D}_{\text{SFT}}$ is the prompts in the SFT dataset.
Rafailov et al. (2023) later introduced the Direct Preference Optimization (DPO) algorithm, which implicitly optimizes the same objective as PPO. DPO directly optimizes the model by a straightforward contrastive loss, boosting the reward of the preferred generation $y_{c}$ and penalizing the one of the non-preferred generation $y_{r}$ from a prompt $x$. DPO is a RL-free approach which has the following loss:
$-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{c}\mid x)}{\pi_{\text{SFT}}(y_{c}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{r}\mid x)}{\pi_{\text{SFT}}(y_{r}\mid x)}\right)$
where $\sigma$ is the sigmoid function, and $\beta$ the scaling parameter. In this study, we lack access to a preference dataset for DPO fine-tuning. Consequently, following the methodology of Rafailov et al. (2023), we generate this dataset by sampling responses from the same SFT dataset, and we rank those preferences using the same reward model (see Appendix A.3). This is justified as it is not possible to obtain a preference dataset from human feedback in the AO setting.
4 Experimental Setup
In this section, we describe the datasets involved for training and evaluation of our resulting models, and present our custom reward targeting the open-world authorship verification and multi-task text embeddings to learn this AO task. We then evaluate the resulting obfuscation against text edition and rewriting baselines.
4.1 Datasets
Training
We use a separate dataset to train our PO models. We fine-tune our base simplification model on the Yelp reviews dataset [^4] https://hf.co/datasets/yelp_review_full Zhang et al. (2015) composed of reviews from Yelp. The dataset is extracted from the Yelp Dataset Challenge 2015. This dataset is employed in an unsupervised way, to ensure we train our models on a large number of authors.
Evaluation
To evaluate our obfuscation models, we use three different datasets. (i) IMDb62555https://hf.co/datasets/tasksource/imdb62, is a subset of the IMDb Authorship Attribution dataset initially presented by Seroussi et al. (2014). It consists of 62 authors with 1,000 texts per author taken from IMDb movie reviews. The utility task associated with this dataset is the review sentiment. For this, we map the movie rating between 0 and 10 associated with each review to a sentiment between positive and negative. A positive review occurs when the review rating is strictly larger than 5. (ii) The Blog Authorship Corpus [^6] https://u.cs.biu.ac.il/~koppel/BlogCorpus.htm dataset Schler et al. (2006) consists of aggregated blog posts from 19,320 bloggers gathered from blogger.com. We pick the list of 13 topics present in the dataset as the utility task. (iii) The Extended-Brennan-Greenstadt [^7] https://hf.co/datasets/tasksource/Drexel-AMT dataset Brennan et al. (2012) is composed of short paragraphs about scholar subjects gathered from 42 different authors from Amazon Mechanical Turk. The utility task of this dataset is indicated by the “background” column, as a binary classification problem.
For all datasets, we create two subsets containing the texts from 10 and 20 authors. For the Blog Authorship Corpus, we select the authors with the highest number of texts. We select the 10 (resp. 20) first authors listed in IMDb62 and Extended-Brennan-Greenstadt. We report summary statistics of each dataset in Table 1 and refer to every dataset as IMDb, BAC, and AMT followed by the number of considered authors. In summary, IMDb has rather long texts, numerous texts per author with a large associated standard deviation. BAC texts are shorter, with a higher number of texts per author compared to IMDb. Finally, for the AMT dataset, the texts are the longest with few variations, and the number of texts per author is the smallest.
4.2 Reward Models
To perform PO, we build a reward model from two different rewards components targeting respectively text semantics and text authorship, aiming to disentangle privacy and utility to control the trade-off.
For utility, we use a pretrained General Text Embeddings (GTE) Li et al. (2023) to represent the reward as a cosine similarity between GTE before and after obfuscation [^8] We use the gte-large-en-v1.5 from sentence-transformers https://hf.co/Alibaba-NLP/gte-large-en-v1.5. Denote as $\mathrm{GTE}(x)$ the embedding vector of size 1024, our utility reward is defined as:
$R_{util}=\mathrm{cossim}(\mathrm{GTE}(x_{ori}),\mathrm{GTE}(x_{obf}))$
For the privacy reward, we use the Learning Universal Authorship Representations model (LUAR), from Rivera-Soto et al. (2021). LUAR’s goal is to transform a given text into a 512 dimensions embeddeding, such that representations of texts by the same author are closer, according to cosine similarity, than those by other authors.
Denote as $\mathrm{LUAR}(x)$ the embedding vector given by the LUAR model, our privacy reward is defined as:
$R_{priv}=1-\mathrm{cossim}(\mathrm{LUAR}(x_{ori}),\mathrm{LUAR}(x_{obf}))$
where $\mathrm{cossim}$ denotes the cosine similarity.
We obtain our final reward by summing the two previous rewards:
$R=R_{util}+R_{priv}$
All implementations details are listed in Appendix A.1.
4.3 Evaluation
Privacy Metrics
The goal for obfuscation is to change the text in order to reduce as much as possible the attacker accuracy. We employ authorship attribution as an evaluation attacker to simulate an attack scenario when the attacker has already access to some sample data of targeted authors to train an attacker classifier. This is a stronger scenario than directly using the reward model as evaluation, since it only assumes one-to-one comparison between texts. For each evaluation dataset, we train a DeBERTa-v3 He et al. (2021) model as an authorship attribution classifier. We split each evaluation dataset in 80%, 10% 10% for training, validation and testing.We measure the accuracy of the attacker model on each test set.
Utility Metrics
We evaluate the utility loss when performing obfuscation similarly to the privacy classifier. For each downstream task dataset, we train a DeBERTa model to quantify utility preservation after text obfuscation. In addition, we also measure the impact on content preservation and soundness (see Appendix B).
Baselines
We use the following baselines:
Original Text
We measure the performance of utility / privacy classifiers when evaluated on original data, the goal of AO would be to decrease the performance of privacy classifiers without decreasing too much the accuracy of utility classifiers.
Synonyms
As a baseline, we perform a naive text edition using synonyms. We use GPTZzz [^9] https://github.com/Declipsonator/GPTZzzs to process original texts, it employs a dictionary of synonyms to replace a given proportion of words with their counterparts. The goal of this baseline is to evaluate the attacker behavior when very small edits are made in the original text.
ALISON
We use ALISON, a recent state-of-the-art text edition AO model leveraging small replacements using a pretrained BERT model. Replacements spans are computed using a threshold on the explanations of an adversarial authorship attribution classifier trained on each evaluation dataset. We train this classifier on each training and validation set before evaluation.
GPT-3.5
Lastly, we include a comparison with GPT3.5 (gpt-3.5-turbo) Ouyang et al. (2022) as a text generation baseline. We use a simple text obfuscation prompt to capture zero-shot capabilities of GPT-3.5 to perform AO. The prompt used can be found in Appendix A.2.
IMDb BAC AMT
Method 10 Authors 20 Authors 10 Authors 20 Authors 10 Authors 20 Authors
Util. $\uparrow$ Attr. $\downarrow$ Util. $\uparrow$ Attr. $\downarrow$ Util. $\uparrow$ Attr. $\downarrow$ Util. $\uparrow$ Attr. $\downarrow$ Util. $\uparrow$ Attr. $\downarrow$ Util. $\uparrow$ Attr. $\downarrow$
Original 73.51 99.78 79.46 99.80 46.73 61.05 53.80 61.14 100 70.37 86.11 42.86
Synonyms 70.38 94.52 76.60 96.08 46.24 59.06 51.20 58.18 91.67 64.81 86.11 36.90
ALISON 61.88 89.59 65.72 91.02 40.70 40.67 41.00 39.22 91.67 70.37 73.33 35.84
GPT-3.5 63.33 66.67 47.37 35.00 37.20 42.73 44.74 31.27 60.00 44.44 61.11 31.14
SFT 64.51 62.50 39.47 80.00 40.41 32.44 40.10 28.28 90.00 26.85 75.00 21.23
TAROT-PPO 63.54 88.89 47.37 71.67 35.38 29.14 42.30 33.62 90.00 35.19 72.22 17.86
TAROT-DPO 57.14 34.74 60.72 17.34 24.57 23.97 28.39 16.42 86.67 22.22 64.18 16.67
Table 2: Evaluation results (Util: classifier accuracy on utility labels, Attr: authorship attribution accuracy) Best values are bolded.
4.4 Realistic Attackers
So far, the considered privacy evaluation adopts the perspective of an attacker who does not have any knowledge about the obfuscation algorithm used. In a more realistic setting, the attacker can likely identify and reproduce the AO model, and perform more advanced attacks by creating adversarial threat models. Following Zhai et al. (2022), we also evaluate our obfuscation models against two enhanced authorship attribution attackers, better suited to simulate real world attack scenarios. We list the different attackers and their specific aspects:
  • Data Augmentation: The attacker is aware of which algorithm was used to perform AO. This knowledge is used to create a new attribution model combining texts from both the original source and obfuscated texts generated from the AO algorithm.
  • Adversarial: The attacker is also aware of the AO algorithm, but use the obfuscated training samples only as a training set for the authorship attribution classifier. Zhai et al. (2022) showed that this attack setting is the one having the higher performances on text edition obfuscations.
For each attack scenario, we train a new authorship attribution classifier using the same parameters (see Appendix A.4 for hyperparameters) and compare the accuracy change from the original attacker.
4.5 Training new utility models with obfuscated texts
We experiment with a second use case to evaluate the downstream utility of obfuscated texts. We use the obfuscated texts of each method as a new training set for our utility classifier. This is useful to evaluate each method capability to generate useful training data that can be further used to train a new classifier on the same utility task.
5 Results
Downstream Effectiveness
In Table 2, we present the accuracy change of privacy and utility classifiers. We observe that both SFT, PPO and DPO reduce the attacker accuracy compared to text edition methods (Synonyms and ALISON). PO helps to learn a good privacy-utility trade-off by largely improving the privacy of obfuscated texts compared to baselines, while preserving similar utility. We observe that DPO consistently outperforms the PPO algorithm on privacy preservation, while using the same base reward model. DPO is also the best-performing privacy preservation over all baselines, with a notable drop of $82,46\%$ on IMDB-20. Note that the utility decrease is larger for the BAC dataset, which could be explained by the number of short texts contained in the dataset, whose edits affect a lot more the end utility. TAROT-DPO also outperforms GPT-3.5 by providing more utility and less attribution on IMDB-20, AMT-10 and AMT-20. The effectiveness of TAROT-PPO lays in its utility preservation capabilities. While not being as private, the utility drop is reduced on nearly each dataset compared to TAROT-DPO.
Adversarial Attackers
Figure 2 highlights the accuracy of adversarial threat models on the IMDb-10 dataset, both attackers reduce the original attacker accuracy drop once retrained using obfuscated texts for the Synonyms approach and Alison. However, text generation methods (GPT-3.5, SFT, TAROT-PPO and TAROT-DPO) are resistant to adversarial threat models, and only GPT-3.5 and TAROT-DPO suffer from a performance decrease from the data augmentation attack. This encourages the path of generation methods as promising obfuscators. Note that this is the first obfuscation approach that is shown to be resistant to threat models. As Zhai et al. (2022) did not include generation models in their study of AO evaluation.
Utility Preservation after Retraining
Figure 3 presents the accuracy of a new utility classifier once trained with obfuscated texts. We observe that the drop in accuracy caused by obfuscation can be compensated by training a new classifier, with an accuracy increase for all methods. Moreover, generation methods are even better candidates for training data, as the final accuracy is higher than the original classifier accuracy. TAROT-PPO and TAROT-DPO are the best-performing approaches on this dataset. This highlights the possibility of creating obfuscation methods that are both preserving privacy and keeping utility for training purposes.
Figure 2: Adversarial training accuracy results on IMDB-10 (lower is better). Generation models are resistant to adversarial training, compared to text edition methods.
Figure 3: Utility classifier accuracy once trained on IMDB-10 obfuscated texts (higher is better). The red line indicates the classifier accuracy when trained and evaluated on original data. The overall utility always increases after training on obfuscated texts, this is key to compensate the utility drop of generation methods.
Method Output
Original I loved the whole story even though it was a tad corny at times . I think great acting and the content of the story kept it going.
Synonyms I loved the quite whole story very even though it was a tad corny at times. I imagine too outstanding playing and the contents of the story kept it sledding.
ALISON I thoroughly enjoyed the entire story even it did have a tad corny at times. I believe the great acting and the story’s content were the main reasons to keep it going.
GPT-3.5 The entirety of the narrative was quite delightful, despite occasional moments of cheesiness. I believe the stellar performances and the substance of the storyline sustained its momentum.
SFT I loved the whole story. It had many good parts and the writing was excellent. I think great acting and the subject matter of the story kept it going.
TAROT-PPO I loved the whole thing. It was a good story and well-written. It also kept me going at times. I think great acting and the content of the story kept me going.
TAROT-DPO I love the whole story. It’s full of action, personality and humour. It keeps me going, though, and the content keeps me going.
Table 3: Obfuscation example from the IMDb dataset.
Qualitative Analysis
We show an obfuscation example in Table 3 for each method. The base Synonyms obfuscation results in awkward phrasing and less natural language, compromising readability. ALISON maintains coherence and clarity with slight formalization (“thoroughly enjoyed” instead of “loved”). GPT-3.5 significantly rephrases the text using sophisticated language. SFT simplifies and shortens the text, retaining clarity but reducing stylistic nuances. TAROT-PPO simplifies further, introducing some repetition, which makes the text less formal but still clear. TAROT-DPO alters the content more significantly, introducing new themes and repetition that can distract from the original meaning. The application of PO assists the text simplification SFT model in making additional modifications to the text. Although these changes in some cases alter the text’s meaning, they preserve its overall utility.
Ablation Study
As a complement, we perform an ablation study of each component of our reward model in Appendix D. It confirms the importance of using a combination of both privacy and utility rewards to learn this trade-off in obfuscation, especially for PPO.
6 Conclusion
We introduced a novel authorship obfuscation framework that focuses on optimizing the privacy-utility trade-off for a specific downstream data usage. We fine-tuned a text simplification model using two policy optimization algorithms to obfuscate the authorship of a given text, while preserving utility for multiple tasks. Our end-models are tuned using two sentence embedding rewards, one for content preservation and one for privacy, resulting in an unsupervised approach made for the open-world authorship setting. The results obtained help to improve the privacy from state-of-the-art AO methods, while preserving task utility. Additionally, we show that generated texts can be used to retrain utility classifier and increase their performances, while limiting the accuracy of more advanced attackers. This calls for more research to design robust evaluation benchmarks for obfuscation systems, to assess and catch failure cases that can map to different real-world scenarios.
7 Limitations
The use of LM as text generators for obfuscation is not without risks, LM are known for their hallucination capabilities, so even if the downstream task is not affected, there is still a possibility that the trained LM generated plausible but false text from the original text. As we did not study the content preservation of resulting texts, we do not emphasize the risk of spread of misinformation or harm that can be generated by our fine-tuned LM.
Another limitation of our approach is that we rely on very small language models (380M parameters for GPT2-medium, our SFT baseline), which benefits from limited memory usage but suffers from a restricted context size for generation. As a result, our method tends to reduce the text length, especially for longer texts. This limitation could be mitigated by increasing the size of the SFT model.
Finally, these methods can be limited when applied to short texts, as the replacements creates significant changes that directly affect the utility task.
8 Ethical considerations
In this work, we present authorship obfuscation methods that are intended for beneficial purposes (learning insights from data while preserving privacy). But we recognize that this task presents some risks of misuse. It can facilitate harmful activities such as posting misinformation, spam, or harmful content, without accountability because of obfuscation. Moreover, these techniques might infringe on intellectual property rights by obscuring the authorship of creative works, depriving creators of their deserved credit. We strongly encourage users to carefully consider these potential dangers before employing such methods.
Appendix A Experimentation Details
A.1 Hardware and code
We conducted all experiments with Nvidia A30 GPU card with 24GB memory and Intel Xeon Gold 5320 CPU. The main libraries used include Pytorch 2.2.2, Huggingface transformers 4.39.3, datasets 2.19.0, tokenizers 0.15.2, trl 0.8.6, evaluate 0.4.1 and sentence-transformers 3.0.0. Due to memory constraints, models are loaded with float16 mixed precision.
Training time for PPO ranges from 15-20 hours, while time for DPO ranges from 6-12 hours. Evaluation time ranges approximately from 19-32 hours.
A.2 GPT-3.5 prompt
In our study, we compare with zero-shot prompting using GPT-3.5, a model with approximately 175 billion parameters. We obfuscate each text on a paragraph level, where the entire text is obfuscated as a unit. We use the following prompt to generate obfuscated texts: "Rewrite the following paragraph so that the author’s style is obfuscated."
A.3 DPO training
While both PPO and DPO algorithms methods aim to optimize a model’s performance based on a reward function, they differ in their approach to policy optimization. PPO uses a surrogate objective function that approximates the true objective function, while DPO directly optimizes the likelihood of generating a response chosen from a preference dataset over another response. This preference dataset is typically collected by having human annotators compare pairs of responses generated by a model and indicate which one is preferred. However, this protocol is impractical for authorship obfuscation because it is difficult to evaluate with human annotations. Therefore, we apply an initial preprocessing step to generate the preference dataset before DPO fine-tuning. We generate preference pairs from SFT outputs, and rank these preferences using the same reward model as PPO. Algorithm 1 outlines our method for creating this preference dataset for DPO. Preliminary experiments showed that removing samples with closely similar authorship rewards accelerates training convergence. So we specify filtering thresholds $\epsilon_{priv}$ and $\epsilon_{util}$. After testing multiple values, we set $\epsilon_{priv}=0.10$ and $\epsilon_{util}=0.05$
Algorithm 1 Preference Dataset Generation
SFT dataset $\mathcal{D}$, privacy threshold $\epsilon_{priv}$, utility threshold $\epsilon_{util}$
prompts = []
chosen = []
rejected = []
for prompt $\in\mathcal{D}$ do
     left, right = generations from the SFT model
     $R_{util-left}$, $R_{priv-left}$ = privacy and utility rewards from the left obfuscation candidate
     $R_{util-right}$, $R_{priv-right}$ = privacy and utility rewards from the right obfuscation candidate
     if $\|R_{priv-right}$ $R_{priv-left}\|>\epsilon_{priv}$ and $\|R_{util-right}$ $R_{util-left}\|<\epsilon_{util}$ then
         if $R_{priv-right}$ > $R_{priv-left}$ then
              prompt.append(prompt)
              chosen.append(right)
              reject.append(left)
         else
              prompt.append(prompt)
              chosen.append(left)
              reject.append(right)
return prompts, chosen, rejected
A.4 Hyperparameters
Table 4 and Table 5 present hyperparameters used for PO algorithms and evaluation classifiers. Due to limited time and computational resources, we are unable to conduct an exhaustive search across all hyperparameters. Instead, we report the best-performing hyperparameters we identified.
TAROT-PPO TAROT-DPO
learning rate 1.47e-5 2.96e-5
batch size 16 32
# epochs 3 3
KL coef / beta 0.2 0.1
top p 1.0 1.0
temperature 1.0 1.0
Table 4: Training hyperparameters for PO algorithms.
DeBERTa (Privacy) DeBERTa (Utility)
learning rate 2e-5 2e-5
batch size 8 8
# epochs 3 3
Table 5: Training hyperparameters for evaluation models.
A.5 Baseline implementation details
Synonyms
We use GPTZzz to process original texts, it employs a dictionary of synonyms to replace a given proportion of words with their counterparts. The goal of this baseline is to evaluate the attacker behavior when very small edits are made in the original text. We use the FinNLP synonym list and ask the algorithm to change up to 90% of words, and 80% of adjectives.
ALISON
We use the author’s code implementation of ALISON, we use the largest edition parameters ($L=250$ and $c=1$) to edit the final text as much as possible.
GPT3.5
We use the gpt-3.5-turbo API endpoint from OpenAI to compute obfuscation, with default temperature, max_tokens and top_p.
Appendix B Content preservation and soundness study
We also study the impact on content preservation when obfuscating the text with generation models, including TAROT. Table 6 presents multiple content preservation metrics on the IMDB-10 dataset. Naturally, text edition methods obtain the best content preservation scores, compared to generation methods. In contrast, generation methods are superior in terms of linguistic acceptability (CoLA), since they generate the complete text as a whole. TAROT-DPO outperforms other methods on this metric.
Rouge-1 Rouge-2 Rouge-L BLEU METEOR BERT Score CoLA
Original - - - - - - 69.31
Synonyms 83.86 68.61 83.68 64.64 92.41 94.61 30.20
ALISON 98.24 97.08 98.19 67.48 97.61 99.01 43.88
GPT-3.5 38.13 11.90 29.15 6.81 33.61 81.81 73.82
SFT 55.69 34.04 43.20 24.06 41.13 85.58 66.66
TAROT-PPO 51.33 29.36 38.67 20.77 37.93 84.50 74.46
TAROT-DPO 42.52 17.27 29.14 10.77 30.04 80.56 81.10
Table 6: Content preservation scores on the IMDB-10 dataset.
Appendix C Complete Evaluation Results
Figure 4 presents the complete evaluation results of adversarial training on all datasets.
Figure 5 presents the complete utility evaluation after retraining on each dataset. The findings presented for IMDb-10 persist for IMDB-20 and AMT-20. We observe a smaller change in utility over the AMT-10 dataset due to the high base accuracy of the original classifier (1.0). However, this result does not hold for the BAC-10 and BAC-20 datasets, which is due to the lack of utility preserved after obfuscation. The blog authorship corpus dataset consists mainly of short texts, making it challenging for rewriting methods to transform the text without significantly affecting utility. This issue persists even after retraining the classifier on the obfuscated data.
Figure 4: Adversarial training accuracy results (lower is better).
Figure 5: Utility classifier accuracy once trained on obfuscated texts (higher is better). The red line indicates the classifier accuracy when trained and evaluated on original data.
Appendix D Reward model ablation study
We perform a reward model ablation study to evaluate the importance of each reward component. Table 7 presents the reward value after training on different setups. We observe that the utility preservation and privacy components are both necessary to balance the privacy-utility trade-off. When we remove the LUAR-based reward, it leads to better GTE similarity at the expense of privacy. Similarly, removing the GTE reward leads to better privacy scores at the expense of utility. In practice, removing the privacy reward leads to models that try to copy the original text. While removing the utility reward leads to very short text, with only few words.
Method TAROT-PPO TAROT-DPO
LUAR GTE LUAR GTE
No privacy 0.975 0.993 0.983 0.977
No utility 0.403 0.421 0.706 0.633
No ablation 0.931 0.825 0.915 0.738
Table 7: Reward model values when removing one component. A high LUAR value indicates low privacy, and a high GTE value high utility.
Appendix E Scientific Artifacts
We list in this section the licenses used in this paper:
Models
DeBERTa-v3 (MIT) Keep It Simple (apache-2.0) LUAR (apache-2.0) GTE (apache-2.0)
Software
GPTZzz (GPL-3.0) ALISON (MIT) GPT-3.5 (Terms of use [^10] https://openai.com/policies) Pytorch (BSD-3) Huggingface transformers, transformers, datasets, trl, evaluate and sentence-transformers (apache-2.0)